What is the osmotic pressure of an aqueous solution containing 0.2 mol of dissolved potassium chloride (completely dissociated) and 2 g of glucose at 37 ° C in 500 cm3?


1
Expert's answer
2021-05-19T02:26:29-0400

См = n в-ва / Vр-ра 

m(KCl)=0.2mol*74.5g/mol=14.9g

V=22.4g/mol*0.2mol=4.48l

n(C6H12O6)=2g/180g/mol=0.0111mol

V(C6H12O6)=0.0111mol*22.4g/mol=0.249l

Cm=(0.0111+0.2)mol/(4.48+0.249)l=

=0.211mol/4,729l=0.045mol/l


P=Cm*R*T

P=0.45*8.314*31=11.6